OpenStudy (anonymous):

The acceleration function and the initial velocity for a particle moving along a line are: a(t)= 2t + 3, v(0)= -4, 0

OpenStudy (anonymous):

You're given a(t) , which is the change of velocity over time. Integrate this function with respect to t, and you have a velocity function. Don't forget that when you integrate there is a constant C, ( this should be the velocity when time equals zero) After you have solved for this constant, integrate again to get a position function, but this time integrate from t= 0 to 3 and you should have your answer

OpenStudy (anonymous):

a(t)= 2t + 3, v(0)= -4, 0<t<3 distence travelled , s(t)= ut + integral . ( 1/2 a(t)^2) = -12+1/2*integral (a(t))^2